OpenStudy (anonymous):

8. Find S8 for the geometric series 3 + -6 + 12 + -24 +… a. -255 b. 768 c. -96 d. 192

OpenStudy (anonymous):

easy use ar0+ar1+ar2 so a is 3

OpenStudy (anonymous):

r is -6=3*r r=-2

OpenStudy (anonymous):

so then 3*(-2)^57

OpenStudy (anonymous):

no wait i thought you meant term lol

OpenStudy (anonymous):

but stil a and r are valid

OpenStudy (anonymous):

\[\sum_{n=0}^73\times (-2)^n=3\times \frac{(-2)^8-1}{-2-1}\]

OpenStudy (anonymous):

use this formula \[{\frac {a{r}^{n+1}}{r-1}}-{\frac {a}{r-1}} \] in it i showed you a and r n is the ith_term-1 which is 58-1 in your case

OpenStudy (anonymous):

what timo86m said is correct, \(a=3,r=-2\) formula to use is \[\sum_{k=0}^nar^k=a\frac{r^{n+1}-1}{r-1}\]

OpenStudy (anonymous):

i get \(-255\)

OpenStudy (anonymous):

\[\sum_{n=0}^73\times (-2)^n=3\times \frac{(-2)^8-1}{-2-1}=3\frac{256-1}{-3}=-255\]